Six jurors are to be selected from a pool of 20 potential candidates to hear a civil case involving a lawsuit between two families. Unknown to the judge or any of the attorneys, 4 of the 20 prospective jurors are potentially prejudiced by being acquainted with one or more of the litigants. They will not disclose this during the jury selection process. If 6 jurors are selected at random from this group of 20, find the probability that the number of potentially prejudiced jurors among the 6 selected jurors is none
Answer #1

Solution. the number of potentially prejudiced jurors among the 6 selected jurors is none P(none) = 4C0×26C6/20c6= 1×8008/38760=0.2266 ,so P(none)=0.2266 thank you

.